OpenStudy (anonymous):

use distributive property to write with no parenthesis then simplify results. 2(3x-7) && -4(x+3)-14

jimthompson5910 (jim_thompson5910):

I'll do the first and it should help you do the second 2(3x-7) 2*3x+2*(-7) ... Distribute 6x - 14 So 2(3x-7) simplifies to 6x - 14

OpenStudy (anonymous):

Thank You!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

just to makw sure i got it...the answer to the second onw is -5x+20?

jimthompson5910 (jim_thompson5910):

no, that's incorrect

jimthompson5910 (jim_thompson5910):

-4(x+3)-14 -4*x-4*3-14 -4x - 12 - 14 Did you get this far?

OpenStudy (anonymous):

sorry looking at the wornf problem.!

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

-4x+26? or do i multiply

jimthompson5910 (jim_thompson5910):

you're close

jimthompson5910 (jim_thompson5910):

but -12 - 14 isn't +26

OpenStudy (anonymous):

2? i get confused when it come to that last part, im not sure whether i multiply or what :/

jimthompson5910 (jim_thompson5910):

in the last part, you combine -12 and -14 to get -26

jimthompson5910 (jim_thompson5910):

so the answer is really -4x-26

OpenStudy (anonymous):

oh so i add and keep the - bc its - on both?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

-12 - 14 = -(12 + 14) = -(26) = -26

OpenStudy (anonymous):

BUT if it were a + && - i would subtract right? so - stays - and + stays +

jimthompson5910 (jim_thompson5910):

Yes, if you had +14 - 12, you subtract if you had -14 + 12, you subtract You're subtracting the numbers (ignoring the negative signs). The result is positive if the larger number is positive. The result is negative if the larger number is negative So 14 -12 = 2 and -14+12 = -2

OpenStudy (anonymous):

ok thank you... hopefully i remember and dont get confused onmy test! thanks!

jimthompson5910 (jim_thompson5910):

you're welcome